A quantity, 2/3 of it, 1/2 of it, and 1/7 of it, added together, equals 3. What is the quantity?
Answer by edjones(8007)

Let the quantity be x
x + 2x/3 + x/2 + x/7 =3
42x+28x+21x+6x=126 multiply each side by 42 to eliminate fractions.
97x=126
x=126/97
Ed